OpenStudy (wesdg1978):

The area of the base B and the height h of a pyramid are related to the pyramid's volume V by the formula V =1/3Bh. How is dV/dt related to dh/dt if B is constant? I chose C, but my teacher said try again? The available answers are: A) dV/dt =B/3(dh/dt) B) dV/dt = B dh/dt C) dV/dt =1/3 (dh/dt) D) dV/dt =dh/dt Would it just be D?

OpenStudy (ybarrap):

V = (1/3) B * h (dV/dt) = (1/3) B * (dh/dt ) + (1/3)*(dB/dt) * h , by chain rule = (1/3) B * (dh/dt)